# ARC: A Generalist Graph Anomaly Detector with In-Context Learning

### v1

@@ F.3 Details of Implementation
Implementation Pipeline. We employ a fixed set of hyper-parameters to build a generalist GAD model for all datasets. First, we train all the methods (including baselines and ARC) on the training set with full labels. Then, the methods are evaluated on each dataset from respectively. For baselines that require fine-tuning, we further conduct dataset-specific tuning at this stage.
@@ G.1 Performance Comparison in Terms of AUPRC
In terms of AUPRC, Table 4 gives comprehensive comparative results with consistent observations with the AUROC results. Specifically, we have the following observations. ❶ ARC still demonstrates strong anomaly detection in generalist GAD scenarios without any fine-tuning. Specifically, ARC achieves state-of-the-art performance on three of the eight datasets and demonstrates competitive performance on the remaining datasets. On several datasets, ARC showed significant improvement over the best baseline (e.g., 121.4% on Cora, 110.2% on Citeseer). ❷ GAD methods that only pre-train specific to a dataset usually result in poor generalization to new datasets. Specifically, existing methods perform very erratically on different datasets, which can be attributed to capturing only specific anomaly patterns. ❸ Using dataset-specific fine-tuning, baseline methods can achieve better performance in most case. However, in some cases the improvement can be small or even negative, demonstrating the limitations of fine-tuning.

### v2

@@ F.3 Details of Implementation
Implementation Pipeline. We employ a fixed set of hyper-parameters to build a generalist GAD model for all datasets. First, we train all the methods (including baselines and ARC) on the training set with full labels. Then, the methods are evaluated on each dataset from respectively. For feature projection, we employ the PCA algorithm to map the raw features into a fixed space with . When the original feature dimension is smaller than the predefined projection dimension , we use a random projection (e.g., Gaussian random projection) to upscale the feature into a higher dimensionality and then unify the dimensions into with the projection strategy. For the baselines that require fine-tuning, we further conduct dataset-specific tuning at this stage.
@@ G.1 Performance Comparison in Terms of AUPRC
In terms of AUPRC, Table 4 gives comprehensive comparative results with consistent observations with the AUROC results. Specifically, we have the following observations. ❶ ARC still demonstrates strong anomaly detection in generalist GAD scenarios without any fine-tuning. Specifically, ARC achieves state-of-the-art performance on five of the eight datasets and demonstrates competitive performance on the remaining datasets. On several datasets, ARC showed significant improvement over the best baseline (e.g., 131.1% on Cora, 98.8% on Citeseer). ❷ GAD methods that only pre-train specific to a dataset usually result in poor generalization to new datasets. Specifically, existing methods perform very erratically on different datasets, which can be attributed to capturing only specific anomaly patterns. ❸ Using dataset-specific fine-tuning, baseline methods can achieve better performance in most case. However, in some cases the improvement can be small or even negative, demonstrating the limitations of fine-tuning.